# Revisiting Few-Shot Object Detection with Vision-Language Models

### v2

@@ Abstract
Few-shot object detection (FSOD) benchmarks have advanced techniques for detecting new categories with limited annotations. Existing benchmarks repurpose well-established datasets like COCO by partitioning categories into base and novel classes for pre-training and fine-tuning respectively. However, these benchmarks do not reflect how FSOD is deployed in practice. Rather than only pre-training on a small number of base categories, we argue that it is more practical to fine-tune a foundation model (e.g., a vision-language model (VLM) pre-trained on web-scale data) for a target domain. Surprisingly, we find that zero-shot inference from VLMs like GroundingDINO significantly outperforms the state-of-the-art (48.3 vs. 33.1 AP) on COCO. However, such zero-shot models can still be misaligned to target concepts of interest. For example, trailers on the web may be different from trailers in the context of autonomous vehicles. In this work, we propose Foundational FSOD, a new benchmark protocol that evaluates detectors pre-trained on any external datasets and fine-tuned on $K$-shots per target class. Further, we note that current FSOD benchmarks are actually federated datasets containing exhaustive annotations for each category on a subset of the data. We leverage this insight to propose simple strategies for fine-tuning VLMs with federated losses. We demonstrate the effectiveness of our approach on LVIS and nuImages, improving over prior work by 5.9 AP.
Figure 1: We propose a new few-shot object detection (FSOD) benchmark which embraces the challenge of aligning multimodal foundation models to target concepts with vision and language. On the left, we describe the standard setup: methods pre-train on base classes (with many examples per class) and then fine-tune on $K$-shots of novel (and optionally base) classes. On the right, we describe our proposed setup: Given the scale and often private nature of data used to train VLMs, it is impractical to maintain a split of base and novel classes. Instead, one should directly fine-tune VLMs on $K$-shots of the target classes (and evaluate only those target classes). Importantly, VLMs allow us to exploit additional language cues such as class names and descriptions for fine-tuning. We show that such “zero-shot” language cues already outperforms state-of-the-art FSOD methods without any fine-tuning.
Figure 2: Poor Alignment Between VLM and Target Concepts. Although VLMs show impressive zero-shot performance, they struggle when the target class is different from concepts encountered in web-scale training. On the left, we see that the nuImages dataset defines the cab of the truck as a separate concept from its trailer (shown in red). In contrast, the VLM predicts the entire vehicle as a truck (shown in green). On the right, we present the actual class definitions given to the nuImages annotators, provided as both textual descriptions and visual examples. Just as human annotators learn concepts from few-shot multi-modal examples, we argue that VLMs should be fine-tuned with $K$ vision-language examples.
@@ 1 Introduction
Object detection is a fundamental problem in computer vision [8, 21] that has matured in recent years [32, 24, 31, 22]. Given a large-scale annotated dataset, one can train a detector from scratch. However, training object detectors for domains with limited annotated data remains challenging, motivating the problem of few-shot object detection.
Status Quo. Few-shot object detection (FSOD) benchmarks have made considerable progress on learning to detect new categories from limited training data. Existing benchmarks are constructed by partitioning popular object detection datasets like PASCAL VOC [5] and COCO [21] into base categories (with many examples per class) and target novel categories (with few examples per class). Detectors are first pre-trained on base classes and are fine-tuned on $K$ examples (or $K$-shots) from novel classes.
Conventional FSOD benchmarks enforce base and novel classes to be disjoint to prevent concept leakage and measure generalization to unseen categories. However, as most detectors are pre-trained on ImageNet, concept leakage already occurs in contemporary benchmarks. For example, cat and person are considered novel in the COCO FSOD benchmark but are already present in ImageNet. Similarly, car is considered novel even though similar concepts like sports car and race car are present in ImageNet. Since concept leakage is difficult to avoid, we argue that it should instead be embraced. Intuitively, pre-training on large-scale diverse base categories (which may overlap with novel concepts) will ultimately improve generalization to novel classes.
Foundational FSOD. Rather than explicitly filtering target classes from pre-training [12, 50, 3], practitioners will likely employ foundational vision-language models (VLMs) pre-trained on (potentially private) web-scale data [33, 14, 20, 35] and fine-tune them for their task. As VLMs’ pre-training datasets contain diverse concepts [30, 20] [^1] The CLIP [30] pre-training dataset contains 500,000 concepts, spanning many categories encountered in the real world., it is challenging to prevent concept leakage. Therefore, one may be hesitant to exploit VLMs for FSOD. However, the performance of foundational models is undeniable; state-of-the-art VLMs like GroundingDINO already dominate all leading FSOD methods on COCO (48.3 vs. 33.1 AP) without fine-tuning (cf. Table 1).
Multi-Modal Concept Alignment. Does the strong zero-shot performance of VLMs imply that few-shot detection is no longer an interesting problem? No! We find that the target class name is often an insufficient description of the target concept. For example, trailer in nuImages [1] is defined differently than trailer in web-scale data. As shown in Fig. 2, even human annotators require few-shot instructions to identify subtle aspects of the target concept [2]. Interestingly, annotator instructions are naturally multi-modal, often including visual examples and textual descriptions. We advocate for a FSOD setup that uses similar visual and language cues for VLM concept alignment. We refer to our proposed setup as Foundational FSOD (cf. Fig. 1).
Federated Few-shot Learning. In order to effectively align VLM concepts with $K$-shot multi-modal “instructions”, we leverage a simple but evidently underappreciated observation: few-shot object detection datasets are actually federated datasets [11]. A federated dataset is a dataset comprised of smaller subsets, where each subset is exhaustively annotated for only a single category. For example, cars may or may not appear in the background of the $K$ images annotated with motorcycles (cf. Fig. 3). However, existing FSOD methods incorrectly assume that no cars are present in the background of non-car images. Inspired by prior work in learning with federated datasets [48] and weakly-supervised learning [15, 40, 39], we demonstrate that fine-tuning VLMs with federated losses consistently improves over zero-shot inference (cf. Tables 2, 4).
Contributions. We present three major contributions.
  1. 1.
@@
    We modernize FSOD benchmarks by embracing vision-language foundation models that are pretrained on internet-scale data. We highlight the practical challenge of using multi-modal few-shot examples to define the target semantic concept (as shown in Fig. 2).
  2. 2.
    We point out that existing FSOD benchmarks are actually federated datasets, and present simple strategies for fine-tuning VLMs.
  3. 3.
    We conduct extensive experiments to ablate our design choices and demonstrate that our simple approach achieves state-of-the-art results on the LVIS and nuImages Foundational FSOD benchmarks.
@@ 2 Related Works
Few-Shot Object Detection aims to detect new categories with limited training data [16]. Recent work explores two primary approaches: meta-learning and transfer learning. Meta-learning-based methods focus on acquiring generalizable features from a set of base classes, which can then be applied to identify novel classes. For example, [13] proposes a technique that re-weights features from base classes to predict novel classes. [43] proposes a framework addressing both few-shot object detection and few-shot viewpoint estimation. [6] introduces a general FSOD network that learns a matching metric between image pairs, while [41] enhances object features using a universal prototype. More recently, [45] proposes a generative approach that is robust to noisy object proposals for novel classes. In contrast, transfer learning involves partially freezing network weights pretrained on a base dataset to improve a model’s ability to generalize to novel classes with limited data. Transfer learning approaches often follow a two-stage fine-tuning strategy: first train on base classes and then fine-tune the box classifier and regressor with K-shots from novel classes. This strategy has historically outperformed meta-learning approaches [38]. Recent work has primarily focused on improving classification performance. [36] utilizes a contrastive proposal encoding loss to encourage instance-level intra-class compactness and inter-class variance. Similarly, [19] applies a class margin loss to balance inter and intra-class margins. Our approach leverages transfer-learning by fine-tuning vision-language models (VLMs) pre-trained on large-scale datasets.
@@
Vision Language Models are trained on a large-scale collection of weakly-supervised image-text pairs collected from the web. These models embed images and text into a shared space, enabling open-vocabulary detection. Early works adapt VLMs for object detection by either distilling the model’s predictions for specific image regions [10, 9] or directly incorporating detection components into frozen [17] or fine-tuned [27, 28, 4] encoders. In contrast, RegionCLIP [47] employs a multi-stage training approach, which involves generating pseudo-labels from captioning data, conducting region-text contrastive pre-training, and fine-tuning on detection data. GLIP [20] uses a single text query for the entire image and frames detection as a phrase grounding problem. Detic [49] addresses long-tail detection performance by leveraging image-level supervision. In the context of open-vocabulary detection, there may be some overlap between categories seen during training and testing. We use the term “zero-shot inference” to signify that a model has never been trained on the target dataset.
Federated Datasets are constructed by combining smaller datasets, each resembling a conventional object detection dataset for a single category [11]. Each of these smaller datasets ensures exhaustive annotations for a specific category. Images within each smaller dataset may overlap, resulting in some images with exhaustive annotations for multiple categories. Importantly, since exhaustive annotations for a particular category are only guaranteed within each small dataset, most images are sparsely annotated. As a result, naively training models with federated datasets leads to much sparser gradients [37]. To address this challenge, CenterNet2 [48] introduced FedLoss, a simple modification of cross-entropy loss which randomly samples a subset of negative categories for each image. We adopt FedLoss for FSOD, achieving consistent improvements over zero-shot inference.
Figure 3: Identifying Negative Pseudo-Labels. On the left, we visualize the standard $K$-shot detection setup, which we argue is actually a federated dataset [11] where one is given multiple mini-datasets of $K$ images. In this case, we visualize two $K=1$ datasets of bus and motorcycle. Importantly each mini-dataset does not provide information about the presence of other objects. Previous FSOD methods apparently ignore this fact, and instead assume the collective set of few-shot images are fully annotated across all object classes. This will likely produce many incorrect negative labels – e.g., all unlabeled cars in the background of the motorcycle mini-dataset will be incorrectly treated as negative cars. We use a ✓  to denote that a given image will be treated as a negative example of a given class by the learner and a ✗  to denote that a given image will be ignored when learning a given class. We color such negative labels as green when correct and red when incorrect. Naive FSOD approaches learn about all classes from all images, which results in many incorrect negative labels (shown in red on the left). Instead, we embrace the partially-labeled nature of the data and exploit tools from weakly-supervised learning, such as the use of psuedo-labels predicted by a teacher. We train (or rather, fine-tune) initial detectors on only the appropriate mini-dataset and use thresholded psuedo-detections (shown as cyan boxes on the right) to find images that can be confidently treated as (pseudo) negatives, which results in much fewer mistakes (shown in red on the right). This in turns produces improved performance. We also attempted to learn from psuedo positive labels, but found these to be less reliable. We include pseudo-code in the supplement.
Weakly Supervised Learning techniques are especially popular when learning in data-constrained settings, leveraging large-scale noisy annotations to improve model performance. Prior works [15] learn from noisy annotations using negative labels for classification. Further, [18, 44] leverage pseudo-labels for self-training. We modify our fine-tuning approach to leverage negatives derived from pseudo-labels to improve FSOD performance.
3 FSOD with Vision-Language Models
@@
As shown in Fig 1, our proposed Foundational FSOD benchmark uses vision-language models (VLMs) pre-trained on diverse, large-scale datasets prior to fine-tuning on $K$ examples for each target class. We contrast our proposed setup with standard benchmarks, demonstrate that FSOD benchmarks are actually federated datasets, and present simple strategies for fine-tuning VLMs below.
@@ 3.1 Foundational FSOD Benchmark
Existing FSOD benchmarks repurpose well-established datasets like PASCAL VOC [5] and COCO [21] by partitioning them into base and novel classes for pre-training and fine-tuning, respectively. For COCO, the 60 categories disjoint with PASCAL VOC are used as base classes and the remaining 20 are used as novel classes [38]. However, this setup is artificial and does not reflect how FSOD is deployed in practice. First, it requires FSOD methods to detect common categories such as car and person by assuming they have only few examples. Importantly, VLMs like GroundingDINO [23] can already detect common categories with high accuracy on COCO without fine-tuning (cf. Table 1). Therefore, we focus on benchmarking Foundational FSOD on more realistic and challenging datasets like LVIS [11] and nuImages [1]. In addition, existing FSOD benchmarks require that datasets are partitioned into base and novel classes, which is not feasible for large-scale (often private) foundational datasets. For example, although CLIP’s [30] model weights are publicly available, its pre-training dataset is not. Instead, FSOD methods should only fine-tune VLMs on $K$-shot annotations for $C$ target classes, and also evaluate performance on these $C$ classes.
@@ 3.2 Few-Shot Multi-Modal Concept Alignment
Although VLMs achieve strong zero-shot performance on common classes, they struggle when the target class is different from vision-language concepts encountered in web-scale training (cf. Fig. 2). For example, nuImages [1] defines trailer as independent from the truck cab. However, Detic jointly detects the truck cab and its trailer together. This fine-grained distinction is provided to human annotators with visual examples and textual descriptions. Similarly, we provide few-shot multi-modal examples to align Detic to each target class.
We start by naively fine-tuning Detic [49] on $K$-shot vision-language examples. We ablate the impact of freezing different parts of Detic in Table 5 and find that freezing the backbone, RPN, and classifier head with CLIP embeddings yields the best performance (cf. Figure 4). However, due to the sparsity of annotations in the few-shot examples, we posit that Detic will receive sparser gradients [37]. Importantly, this degrades detection accuracy because all unannotated objects are treated as negatives. Therefore, we explore three strategies for handling negatives as described below.
3.3 FSOD Benchmarks are Federated Datasets
Prior works follow the $K$-shot dataset creation process established by [38]. To construct a $K$-shot dataset, we select a target class $C$ and an image at random. If the total annotations for class $C$ in the image are less than or equal to $K$, we add the image to our dataset. We repeat this process for all classes until we have exactly $K$ annotations per class. Importantly, each image in the dataset is exhaustively annotated for a subset of all classes. Recall, a federated dataset is also comprised of images that are exhaustively annotated for a specific category. This suggests that we can leverage insights about federated datasets [11, 48] to train better few-shot object detectors.
Fine-Tuning with FedLoss. We fine-tune Detic with Federated Loss (FedLoss) [48] using a subset $S$ of classes $C$ for each training image. Specifically, we use a binary cross-entropy loss on all classes in $S$ and ignore classes outside of $S$ during training. $S$ is comprised of the ground-truth annotation class along with randomly sampled negative classes for each image. We sample these negative classes in proportion to their square-root frequency in the training set. We find that probablistically sampling negatives rather than labeling all unannotated classes as negatives improves fine-tuning results, reliably beating zero-shot performance. Importantly, although FedLoss has been explored in the context of long-tailed detection, applying it to FSOD provides considerable performance improvements, reaffirming that FSOD benchmarks are actually federated datasets.
Fine-Tuning with Inverse FedLoss. FedLoss samples common classes like car more frequently as negatives, hurting detection accuracy for long-tailed datasets like LVIS and nuImages. Instead, we propose Inverse FedLoss (InvFedLoss), a minor modification of FedLoss that samples negative categories in proportion to the inverse of their square root frequency. This ensures that we sample rare categories as negatives more frequently to better match the true data distribution. Leveraging this insight improves over FedLoss and naive fine-tuning.
Fine-Tuning with Pseudo-Negatives. Despite the effectiveness of InvFedLoss, probablistically sampling negatives using dataset-wide statistics is sub-optimal because it does not consider the content of each image. We can improve the accuracy of sampled negatives with pseudo-labels to determine which classes are likely not in a particular image. If the maximal score for any class prediction is less than a threshold, we consider this class to be a negative. Using image predictions to identify pseudo-negatives yields better results than simply using dataset-wide statistics. We find that this strategy works the best. We present pseudo-code in the supplement.

### v3

# references.bib

@@ Abstract
The era of vision-language models (VLMs) trained on large web-scale datasets challenges conventional formulations of “open-world" perception. In this work, we revisit the task of few-shot object detection (FSOD) in the context of recent foundational VLMs. First, we point out that zero-shot VLMs such as GroundingDINO significantly outperform state-of-the-art few-shot detectors (48 vs. 33 AP) on COCO. Despite their strong zero-shot performance, such foundational models may still be sub-optimal. For example, trucks on the web may be defined differently from trucks for a target application such as autonomous vehicle perception. We argue that the task of few-shot recognition can be reformulated as aligning foundation models to target concepts using a few examples. Interestingly, such examples can be multi-modal, using both text and visual cues, mimicking instructions that are often given to human annotators when defining a target concept of interest. Concretely, we propose Foundational FSOD, a new benchmark protocol that evaluates detectors pre-trained on any external datasets and fine-tuned on multi-modal (text and visual) $K$-shot examples per target class. We repurpose nuImages for Foundational FSOD, benchmark several popular open-source VLMs, and provide an empirical analysis of state-of-the-art methods. Lastly, we discuss our recent CVPR 2024 Foundational FSOD competition and share insights from the community. Notably, the winning team significantly outperforms our baseline by 23.9 mAP!
@@ 1 Introduction
Vision-language models (VLMs) trained on (often proprietary) web-scale datasets have disrupted traditional notions of the "open-world," particularly for few-shot recognition. In this paper, we revisit few-shot object detection (FSOD) in the context of these foundational models, propose a new benchmark protocol that allow foundational models to “enter the conversation", and present several simple baselines.
First, we highlight that zero-shot VLMs like GroundingDINO demonstrate a remarkable improvement over state-of-the-art few-shot detectors (48.3 vs. 33.1 AP) on COCO, as shown in Table 1. In hindsight, this is not surprising, as the former is pre-trained on far more data (that may include visual examples of the target concept), while the later is pre-trained on data that is explicitly curated to avoid target concepts of interest. From this perspective, VLMs violate the current training protocol of few-shot benchmarks, suggesting that such protocols need to be rethought in the foundational era.
Concept Alignment. Despite their impressive performance, foundation models used in a zero-shot fashion can still be sub-optimal. For example, trucks as defined for a particular target application like perception for autonomous vehicles may differ from trucks as found on the web (cf. Fig. 1). Indeed, this well-known observation has created the ad-hoc practice of prompt engineering, where users actively search for a textual prompt that elicits the desired zero-shot behaviour. Instead, we argue that one can principally address the challenge of aligning foundation models to target concepts through the lens of few-shot recognition, by presenting VLMs with a few examples of the target concept. Crucially, such examples can be multi-modal, using both text and visual cues, mimicking the natural few-shot multi-modal instructions that are often given to human annotators when defining a target concept of interest [chang2023thinking]. Before introducing our new protocol, we first review the conventional FSOD setup below.
Figure 1: Poor Alignment Between Vision Language Models (VLMs) and Target Concepts. Although VLMs show impressive zero-shot performance, they struggle when the target class is different from concepts encountered in web-scale training. On the left, we see that the nuImages dataset [caesar2020nuscenes] defines the cab of the truck as a separate concept from its trailer (shown in red). In contrast, the VLM predicts the entire vehicle as a truck (shown in green). Similarly, nuImages annotations dictate that a person riding a bicycle must also be labeled as part of bicycle (shown in red) unlike the VLM prediction (in green). On the right, we present the actual class definitions given to the nuImages annotators, provided as both textual descriptions and visual examples. Just as human annotators learn concepts from few-shot multi-modal examples, we argue that VLMs should be aligned to $K$ vision-language examples.
Conventional FSOD. Existing FSOD benchmarks partition object detection datasets like PASCAL VOC [Everingham10] and COCO [Lin2014MicrosoftCC] into base and novel classes. Detectors pre-train on base and then learn novel classes given $K$ examples (or $K$-shots). Current protocols enforce base and novel to be disjoint to prevent concept leakage, allowing one to evaluate generalization to the “open-world". However, as most detectors are pre-trained on ImageNet, we point out that concept leakage already occurs in current FSOD protocols. For example, cat and person are deemed novel for COCO-FSOD but are present in ImageNet data used to pre-train detectors [wang2020frustratingly]. Moreoever, car is deemed novel, but similar concepts like sports car and race car are present in ImageNet, illustrating the difficulty of even defining leakage.
Foundational FSOD. We believe that concept leakage should be embraced. Our Foundational FSOD protocol replaces the base pre-training stage with web-scale pre-training, where such data may be proprietary and not fully disclosed [radford2021learning]. We argue that pre-training on large-scale data will be the key enabler for generalization to the open-world. Note that this hypothesis is difficult to even evaluate under current few-shot protocols, motivating our setup. Moreover, another key property is that $K$-shots may include multi-modal examples spanning both images and text, motivating a multi-modal adaptation stage that aligns the VLM to the target concepts (cf. Fig. 2). We repurpose nuImages for our Foundational FSOD benchmark, a challenging dataset due to open-world categories such as debris and pushable-pullable, which also provides multi-modal annotation instructions.
@@
Contributions. We present three major contributions. First, we modernize FSOD benchmarks by embracing vision-language foundation models that are pretrained on internet-scale data. We highlight the practical challenge of using multi-modal few-shot examples to define the target semantic concept (as shown in Fig. 1). Next, we adapt nuImages for Foundational FSOD, evaluate various popular open-source VLMs, and present an empirical analysis of leading methods. Lastly, we highlight the results from our recent CVPR 2024 challenge hosted in conjunction with the Workshop on Visual Perception via Learning in An Open World.
Figure 2: Foundational Few-shot Object Detection (FSOD). Conventional FSOD protocols (left) allow for pre-training on base classes (with many examples per class) and then fine-tuning on $K$-shots of novel classes, where novel and base are designed to be disjoint. However, we point out that pre-training datasets such as ImageNet often contain classes similar to the novel classes, highlighting the issue of concept leakage. As preventing concept leakage is difficult (if not impossible) and appears to be artificial in the foundational era, we propose the setup of foundational FSOD (right). Our setup allows for pre-training on massively-large (and potentially proprietary) datasets, typical for foundational VLM models. Since these models can process both text and images, one can utilize such multi-modal $K$-shot examples to align VLMs with the target concepts of interest.
@@ 2 Related Works
Few-Shot Object Detection aims to detect new categories with limited training data [kohler2021few]. Recent work explores two primary approaches: meta-learning and transfer learning. Meta-learning-based methods focus on acquiring generalizable features from a set of base classes, which can then be applied to identify novel classes. For example, [kang2019few] proposes a technique that re-weights features from base classes to predict novel classes. [xiao2022few] proposes a framework addressing both few-shot object detection and few-shot viewpoint estimation. [fan2020few] introduces a general FSOD network that learns a matching metric between image pairs, while [wu2021universal] enhances object features using a universal prototype. More recently, [xu2023generating] proposes a generative approach that is robust to noisy object proposals for novel classes. In contrast, transfer learning involves partially freezing network weights pretrained on a base dataset to improve a model’s ability to generalize to novel classes with limited data. Transfer learning approaches often follow a two-stage fine-tuning strategy: first train on base classes and then fine-tune the box classifier and regressor with $K$-shots from novel classes. This strategy has historically outperformed meta-learning approaches [wang2020frustratingly]. Recent work has primarily focused on improving classification performance. [sun2021fsce] utilizes a contrastive proposal encoding loss to encourage instance-level intra-class compactness and inter-class variance. Similarly, [li2021beyond] applies a class margin loss to balance inter and intra-class margins. Our approach leverages transfer-learning by fine-tuning vision-language models (VLMs) pre-trained on large-scale datasets.
@@
Vision Language Models are trained on a large-scale collection of weakly-supervised image-text pairs collected from the web. These models embed images and text into a shared space, enabling open-vocabulary detection. Early works adapt VLMs for object detection by either distilling the model’s predictions for specific image regions [gu2021vild, gu2021open] or directly incorporating detection components into frozen [kuo2022fvlm] or fine-tuned [minderer2022owlvit, minderer2023owlvit2, du2022learning] encoders. In contrast, RegionCLIP [zhong2022regionclip] employs a multi-stage training approach, which involves generating pseudo-labels from captioning data, conducting region-text contrastive pre-training, and fine-tuning on detection data. GLIP [li2021grounded] uses a single text query for the entire image and frames detection as a phrase grounding problem. Detic [zhou2022detecting] addresses long-tail detection performance by leveraging image-level supervision. In the context of open-vocabulary detection, there may be some overlap between categories seen during training and testing. We use the term “zero-shot inference” to signify that a model has never been trained on the target dataset.
Fine-Tuning Foundation Models is of significant interest across many application areas [hu2021lora, zhang2023adding, gao2024clip]. Standard fine-tuning procedures employ both linear probing [chen2020simple, he2022masked, he2020momentum] and full-finetuning [wang2017growing, wu2023revisiting, kirkpatrick2017overcoming] to adapt models to downstream tasks. However, such methods may be suboptimal as they can be computationally expensive. Instead, recent works like CLIP-Adapter [gao2024clip] and Tip-Adapter [zhang2021tip] fine-tune CLIP using parameter-efficient methods [houlsby2019parameter, zhang2020side, jia2022visual] which optimize lightweight MLPs while keeping the encoder frozen. Similarly, inspired by the success of prefix-tuning in language models [deng2022rlprompt, jiang2020can, haviv2021bertese, gao2020making], prompt adaptation [lu2022prompt, zhu2023prompt, xing2023dual, zhou2022conditional] replaces hand-crafted prompts like "a photo of a {cls}" with learned word embeddings. CoOp [zhou2022learning] and other prompting methods [lu2022prompt, zhu2023prompt, zhou2022conditional] finetune CLIP via prefix-tuning. Different from most prior work, we investigate fine-tuning strategies for VLM-based detectors using few-shot multi-modal examples.
3 Foundational FSOD Benchmark
@@
As shown in Fig 2, our proposed Foundational FSOD benchmark utilizes vision-language models (VLMs) pre-trained on diverse, large-scale datasets, which are then aligned to $K$ examples of each target class. We contrast our proposed setup with standard benchmarks and present simple strategies for fine-tuning VLMs below.
@@ 3.1 Foundational FSOD Benchmark
Existing FSOD benchmarks repurpose well-established datasets like PASCAL VOC [Everingham10] and COCO [Lin2014MicrosoftCC] by partitioning them into base and novel classes for pre-training and fine-tuning, respectively. For COCO, the 60 categories disjoint with PASCAL VOC are used as base classes and the remaining 20 are used as novel classes [wang2020frustratingly]. However, this setup is artificial and does not reflect how FSOD is deployed in practice. First, the FSOD benchmarks construct novel classes by including common concepts such as car and person, and require FSOD methods to detect these common classes by assuming they have only few examples. Importantly, VLMs like GroundingDINO [liu2023grounding] can already detect common categories with high accuracy on COCO without fine-tuning (cf. Table 1). Therefore, we focus on benchmarking Foundational FSOD on more realistic and challenging datasets like nuImages [caesar2020nuscenes]. Second, existing FSOD benchmarks require that datasets are partitioned into base and novel classes, which is infeasible for large-scale (often private) foundational datasets. For example, although CLIP’s [radford2021learning] model weights are publicly available, its pre-training dataset is not. Instead, FSOD methods should only fine-tune VLMs on $K$-shot annotations for $C$ target classes (or novel, as termed in the conventional FSOD benchmark), and also evaluate performance on these $C$ classes.
@@ 3.2 Few-Shot Multi-Modal Concept Alignment
Although VLMs achieve strong zero-shot performance on common classes, they struggle when the target class is different from concepts encountered on the web (cf. Fig. 1). For example, nuImages [caesar2020nuscenes] defines the cab of a truck as a separate concept from its trailer. However, Detic detects the entire vehicle as truck. This fine-grained distinction is provided to human annotators with visual examples and textual descriptions. We explore seven methods for alignment (either explicitly by updating model weights via fine-tuning or in context via prompting) VLMs below.
Prompt Engineering uses expressive descriptions in the text prompt, adding attributes, synonyms or language context, to manually improve the alignment of foundation model outputs to target concepts of interest. In our case, we prompt VLMs with synonyms of the nuImages classes to improve detection accuracy. For example, we augment the language query for pushable-pullable with synonyms like cart and wheel barrow.
Standard Fine-Tuning updates the last few layers of a model to adapt to new target classes. For two-stage object detectors, this typically requires training the box regression and classifier head. However, we find that standard fine-tuning is sub-optimal, motivating our proposed approach below.
Federated Fine-Tuning leverages a simple but evidently underappreciated observation: few-shot object detection datasets are actually federated datasets [gupta2019lvis]. A federated dataset is comprised of smaller mini-datasets, where each mini-dataset is exhaustively annotated for only a single category. For example, cars may or may not appear in the background of the $K$ images annotated with motorcycles. However, existing FSOD methods incorrectly assume that no cars are present in the background of non-car images. We devise a simple loss that incorporates this insight, discussed further in the supplement.
Language Prompt Tuning is an established parameter-efficient strategy [shin2020autoprompt, lester2021power] for updating text embeddings with few-shot examples via fine-tuning. Concretely, for a given language query (e.g. stroller), we first extract a text embedding $P^{0}$ and only fine-tune the text embedding [li2021grounded].
Visual Prompting uses images of target concepts that are difficult to describe through text as prompts to learn novel concepts in-context. For example, although debris is a difficult catchall category to define through text, we can use image examples to improve concept alignment. Typically, visual prompts are tokenized and fed as inputs to a frozen VLM.
Multi-Modal Prompting combines language and visual prompting to leverage multi-modal features. Intuitively, multi-modal cues can yield better alignment than uni-modal cues alone; in the above case, ambiguous concepts such as debris can be clarified with both textual descriptions (e.g trash can and tree branch) and visual examples (similar to the multi-modal annotator instructions in Fig. 1!). Here, visual and language prompts are tokenized and separately fed as inputs to a frozen VLM. Specifically, MQDet[xu2024multi] introduces a lightweight module: Gated Class Scalable Perceiver, that fuses visual cues and text descriptions in the text encoder via class-wise cross attention layers.
Multi-Modal Chat Assistants can accomplish many of the same tasks as multi-modal prompting through a multi-modal turn-by-turn conversational interface. However, unlike multi-modal prompting, conversational interfaces allow users to iteratively refine concept definitions, similar to how human annotators often require several rounds of feedback to fully understand the target concept.
4 Experiments
We conduct extensive experiments to validate that zero-shot inference from VLMs significantly improves over state-of-the-art FSOD approaches, suggesting that existing benchmarks should be re-framed to allow foundation models to “enter the conversation”. Moreover, we demonstrate that leveraging language cues, especially those available for free (e.g. class names), are crucial to improving performance on data-constrained tasks like FSOD.
Datasets and Metrics. We repurpose nuImages [caesar2020nuscenes] to support the study of Foundational FSOD. This dataset annotates 18 classes, which are divided into groups with many, medium, and few examples [peri2023towards, ma2023longtailed]. We report AP for each cohort. Although this dataset is not traditionally used for FSOD, nuImages’ open-world categories like debris and pushable-pullable make it particularly challenging (even for VLMs), and is a realistic benchmark for Foundational FSOD. We follow the $K$-shot dataset creation process established by [wang2020frustratingly], described below. To construct a $K$-shot dataset, we select a target class $c$ and an image at random. If the total annotations for class $c$ in the image are less than or equal to $K$, we add the image to our dataset. We repeat this process for all classes until we have exactly $K$ annotations per class. Since the specific $K$ examples can have a significant impact on the overall performance, we run each experiment over 3 random data splits and report the average.